HUBBLE SPACE TELESCOPE DAILY REPORT #2248 PERIOD COVERED: 0000Z (UTC) 11/10/98 - 0000Z (UTC) 11/11/98 Daily Status Report as of 315/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Five Sets of NIC/2 7233 (Dust Disks Around Main Sequence Stars) The Near Infrared Camera (Camera Two) was used to make observations of the dust disk around main sequence stars. The observations were made with the coronagraph to minimize the effects of glare from the bright central star. A dark calibration image was also taken. As described in yesterday's report and HSTAR 6767, the acquisition and re-acquisition for the fourth iteration of this proposal failed, and the take data flag remained down, affecting five observations. Otherwise, the observations were completed as scheduled, and no further anomalies were reported. 1.3 Completed NIC/1 7845 (The Superluminal Source GRS 1915+105) The Near Infrared Camera (Camera One) was used to measure the proper motion of the center of mass of GRS 1915+105 to improve our understanding of the genesis and fate of black holes. GRS 1915+105 is a well-established black hole binary system with superluminal ejections. The proposal completed as planned, with no reported anomalies. 1.4 Completed Seven Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.5 Completed STIS/CCD/MA1 7299 (Interfaces Between Hot and Cold Gas: The Superbubble N51D in the LMC) The Space Telescope Imaging Telescope (CCD/MA1) was used to take E140M echelle spectra for four probes of the X -ray bright superbubble SK-67.104 in the LMC, and then to compare those spectra with WFPC2 images of the H Alpha and S II emission in this region to be taken in parallel. The observation completed without incident. 1.6 Completed Two Sets of WF/PC-2 7619 (WF/PC-2 Cycle 7 Decontamination) The WF/PC-2 was used to conduct a Cycle 7 decontamination procedure. Ultraviolet-blocking contaminants were removed, and hot pixels cured, by warming the CCDs to +20C for 6 hours. A variety of internal images were taken before and after the decontamination in order to verify its success and to monitor the instrument health. The procedures were conducted as scheduled, and no problems were reported. 1.7 Completed Four Sets of NIC/1/2/3 7961 (Flats: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. These flats will provide a monitor for particulate comtamination and detector lateral position {from the coronagraphic spot and FDA vignetting}. The proposal completed with no reported problems. 1.8 Completed Six Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 7623 (WF/PC-2 Cycle 7 Internal Flats) The WF/PC-2 was to be used to make a set of internal Cycle 7 calibration flats which are designed to monitor the pixel to pixel flat field response and the visflat lamp degradation, as well as detect any possible changes due to contamination. As described in yesterday's report and HSTAR 6767, the acquisition and re-acquisition for the fourth iteration of this proposal failed, and the take data flag remained down. 1.10 Completed Two Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. As described in yesterday's report and HSTAR 6767, the acquisition and re-acquisition for the first iteration of this proposal failed, and the take data flag remained down. Otherwise, the observations were completed as planned, and no further anomalies were noted. 1.11 Completed WF/PC-2 6557 (WF/PC-2 Imaging of Intermediate Redshift Absorption Selected Galaxies) The WF/PC-2 was used to make observations of the galaxy Q0058-FIEL where a halo appears to be absorbing light from the background quasi-stellar object. The observations were completed as planned, and no problems were reported. 1.12 Completed WF/PC-2 6838 (Dynamical Properties of Distant Field Galaxies) The WF/PC-2 was used to make observations of distant galaxies contained in a location known as Field-0017. A sample of several hundred field galaxies to I = 24 (c.f. toI=22 z~0.5 from Lilly et al. 1995) is being observed with the Keck telescope in a major new survey. This HST observational series is designed to match the field-of-view of the Keck multiple-object spectrograph. The observations were completed as scheduled, and no problems were noted. 1.13 Completed Three Sets of NIC/2 7329 (The Nature of the Damped LyAlpha Absorbers -- A New Study of Young Galaxies) The Near Infrared Camera (Camera Two) was used to make observations of the young galaxies Q2239-3826, Q1151+0651 and Q2206-1958 in order to study the nature of its damped Lyman-Alpha Absorbers (DLAs). The DLA line systems, occasionally found in quasar spectra, provide one of only a few practical methods of detecting and studying significant numbers of galaxies at high redshifts. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed Two Sets of NIC/1/2/3 7886 (NICMOS Snap Shot Survey of Early-Type Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the early-type galaxies NGC-2907 and NGC-1600 as part of a snap-shot broad-band imaging survey. Since almost all early-type galaxies harbor dust, these images which are less sensitive to extinction and will provide superior light profiles for dynamical studies of the cores of these galaxies. The observations were completed as scheduled, and no problems were reported. 1.15 Completed NIC/2 7227 (A Search for Low Mass/Sub-Luminous Companions to M-Stars) The Near Infrared Camera (Camera Two) was used to make a search for low mass/sub-luminous companions to low mass M class stars. Knowledge of stellar and sub-stellar masses and luminosities at and below the ~0.08 M{sun} H-burning limit is of fundamental importance. A dark calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.16 Completed Two Sets of NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the stars HD180617 and HD239960. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.17 Completed NIC/1 7892 (Narrow-band Imaging of BAL QSOs) The Near Infrared Camera (Camera One) was used to make narrow band observations of the BL-Lac (BAL) QSO Q1333+2840. The detection of extended emission-line gas structures in BAL QSOs would offer another means of investigating the BAL QSO phenomenon as well as establishing a new probe of material in galaxies at high redshift. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 18 Successful : 18 Scheduled Reacquisitions : 3 Successful : 3 2.2 FHST Updates: Scheduled : 36 Successful : 36 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared three times. The NICMOS error counter was cleared at 314/1122Z per ROP NS-11. A TTR was written to document IP data hits at 314/1410Z, at 314/1620Z, at 314/1640Z, and at 314/2140Z. The RMGA was powered on at 314/2105Z and then powered off at 314/2301Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH